OpenStudy (anonymous):

Use the exact values of the sin 30˚, sin 45˚, cos 30˚, and cos 45˚ to find the exact value of cos 15˚.

jimthompson5910 (jim_thompson5910):

because 15 = 45 - 30, we can say cos(15) = cos(45 - 30) now use the identity cos(x-y) = cos(x)*cos(y) + sin(x)*sin(y)

jimthompson5910 (jim_thompson5910):

in this case x = 45 y = 30

OpenStudy (anonymous):



OpenStudy (anonymous):

is that it

jimthompson5910 (jim_thompson5910):

what is cos(45) use the unit circle to determine this

OpenStudy (anonymous):

i have no idea

jimthompson5910 (jim_thompson5910):

do you have a unit circle?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

is it radical 2/2

jimthompson5910 (jim_thompson5910):

yes, \[\Large \frac{\sqrt{2}}{2}\]

jimthompson5910 (jim_thompson5910):

what is cos(30) ?

OpenStudy (anonymous):

radical 3/2

OpenStudy (anonymous):

sin 45= √2/2

jimthompson5910 (jim_thompson5910):

good, sin(30) = ??

OpenStudy (anonymous):

sin 30=1/2

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

good \[\Large \cos(15) = \cos(45 - 30)\] \[\Large \cos(15) = \cos(45)\cos(30) + \sin(45)\sin(30)\] \[\Large \cos(15) = \frac{\sqrt{2}}{2}*\frac{\sqrt{3}}{2}+\frac{\sqrt{2}}{2}*\frac{1}{2}\] \[\Large \cos(15) = \frac{\sqrt{2}*\sqrt{3}}{2*2}+\frac{\sqrt{2}*1}{2*2}\] \[\Large \cos(15) = \frac{\sqrt{2*3}}{2*2}+\frac{\sqrt{2}}{2*2}\] \[\Large \cos(15) = \frac{\sqrt{6}}{4}+\frac{\sqrt{2}}{4}\]

jimthompson5910 (jim_thompson5910):

actually, I just noticed that you had 2's and not 4's so you were close

OpenStudy (anonymous):

understood

OpenStudy (anonymous):

since you are multiplying fraction you have to multiply both the numerator and the denominator

jimthompson5910 (jim_thompson5910):

exactly, multiply straight across

OpenStudy (anonymous):

ok thank you very much i hope to get help from you next time